# The Coral Protocol Open Infrastructure Connecting The Internet of Agents

### v1

@@ 4   The Coral Ecosystem
The Coral ecosystem is organized into three major contexts: AI Agent Developers and Users, Coralized AI Agents, and the Coral Protocol (see Figure 1). These contexts encapsulate the roles, components, and interactions that enable Coral’s secure and interoperable multi-agent environment. In this section, we describe each context and its components in detail, and explain the data flow and interactions between them in the overall architecture.
@@ 4.1   AI Agent Developers and Users
This context represents the human actors (developers and end-users) and their toolchain for interfacing with the Coral ecosystem. Developers contribute to Coral by integrating external AI capabilities and resources. They use specialized Coralizer modules to onboard different assets into the ecosystem: the MCP Coralizer connects external model endpoints via the Model Context Protocol (MCP), allowing external AI models or services to communicate in Coral’s standard format; the Data Coralizer links external data sources (e.g., databases, knowledge bases or live data streams) into the Coral ecosystem, making those data accessible to AI agents; and the Agent Coralizer wraps existing AI agents or services (such as pre-existing language models or automation scripts) to comply with Coral’s protocols and interfaces. Through these coralizers, developers effectively “Coralize” their models, data, or agents — registering them into the ecosystem so that they become available as Coralized AI Agents in the middle context. In addition to these, developers can utilize Other Tools (e.g., software development kits, testing frameworks, or monitoring dashboards provided by the Coral platform) to build, debug, and optimize their agents and integrations.
@@
End-users in this context are the consumers of the AI agent services. They interact with Coralized AI Agents by issuing natural language queries or commands through user-facing applications or interfaces built on top of the Coral Protocol. Users need not be aware of the underlying complexity; they simply pose questions or tasks in everyday language. For instance, a user might ask, “What does AgentX think about hypothesisY?” or instruct the system with a command like, “Pay @AgentZ \$50 to investigate the total area of region R.” These inputs from users enter the Coral ecosystem via the Interaction Mediation service of the Coral Protocol (described later), which ensures the queries are routed appropriately. In summary, the Developers and Users context supplies the ecosystem with integrated AI capabilities (via coralizers) and the driving queries or tasks (from users) that initiate agent interactions.
4.2   Coralized AI Agents
@@
The middle tier of the architecture consists of the Coralized AI Agents — the ensemble of AI services and agents that have been onboarded into the ecosystem and operate under Coral’s framework. Each Coralized agent is an AI service (often backed by a large language model or specialized AI module) that adheres to the Coral Protocol for communication and security. Once developers coralize a model, data source, or existing agent, it becomes a Coralized AI Agent accessible in this layer. Collectively, these agents form a distributed, interoperable multi-agent system, each capable of understanding and responding to natural language prompts.
@@
When an end-user query or command enters the system (for example, the questions or instructions mentioned above), the Coral Protocol’s Interaction Mediation component will dispatch it to the appropriate agent or agents in this layer. An agent specialized in a certain domain (say, Agent X in the earlier query) will receive the question directed at it and generate a response drawing on its knowledge or data (which might have been linked via a Data Coralizer). Coralized agents can also initiate interactions with one another under guidance of the Coral Protocol. For instance, if a user’s request is complex and requires multiple skills, the Coral Protocol can intelligently form a dedicated team by discovering and composing the necessary AI agent together. All such inter-agent dialogues are mediated and logged by the protocol to maintain coherence and security.
@@
Crucially, Coralized AI Agents can handle not only informational queries but also procedural commands that involve actions in the ecosystem. For example, when a user says “Pay @AgentZ \$50 to perform task T,” the addressed agent (Agent Z) will be invoked to execute the task T, and the Coral Protocol will engage its secure payment service to transfer the specified amount. Throughout their operation, coralized agents remain decoupled from any particular user interface or platform — they rely on the Coral Protocol to handle incoming requests, outgoing replies, and any coordination with other services. This design allows a heterogeneous collection of AI agents to work together seamlessly: each agent focuses on its specialized processing (e.g., analyzing data, answering questions, performing computations), while the protocol manages communications and shared context. The result is a flexible multi-agent ecosystem where, from the user’s perspective, complex workflows (potentially involving several agents and data sources) can be invoked with simple natural language requests.
@@ 4.3   Coral Protocol
The Coral Protocol is the underlying framework that connects users, developers, and coralized agents, providing the services and infrastructure necessary for secure, coordinated interactions. It comprises two main parts: a set of core Coral Services that mediate and secure the ecosystem’s operations, and the Deployment and infrastructure layer that supports these services. An important additional concept in this context is Secure Team Formation, which refers to the protocol’s capability to dynamically assemble multiple agents into collaborative task teams in a secure manner. We describe each of these aspects below.
@@ 4.3.1 Coral Services
The Coral Protocol offers several key services to orchestrate interactions:
@@
    Secure Multi-Agent Workflow coordinates scenarios where multiple AI agents collaborate on a shared objective. This service sets up and manages the workflow of information and control between agents, enforcing security policies such as authentication of each agent’s identity and permissions. It ensures that agents exchange data through approved channels and that the workflow follows any constraints (for instance, not revealing sensitive data to an agent lacking clearance). Together, these Coral services form the intelligent control plane of the ecosystem, mediating every interaction to guarantee it is executed efficiently and safely.
@@ 4.3.2 Deployment Infrastructure
To support the above services and the execution of AI agents, the Coral Protocol leverages a distributed deployment infrastructure. A network of MCP Servers hosts the AI models and agent runtimes. “MCP” refers to the Model Context Protocol – a standardized interface that these servers implement to allow agents to request model inference or tool usage with a unified API. By deploying agents on MCP servers, Coral ensures that each agent can be invoked remotely with proper context and resource allocation, abstracting away the hardware or environment details. These servers handle the heavy computation of AI tasks (such as running large language model inferences or data processing jobs) and expose endpoints for the Interaction Mediation service to call. In parallel, the Coral ecosystem integrates a blockchain network into its infrastructure. The blockchain serves as a secure ledger for recording important events and transactions: it records payment transactions issued via the Secure Payments service, and can also log multi-agent agreements or task completion records for auditability. By using blockchain technology, Coral introduces an immutable and transparent layer of trust – for example, users and developers can verify that an agent was indeed paid for its service, or that a particular multi-agent task’s outcome was agreed upon by all parties. The blockchain also aids in decentralized governance of the ecosystem, ensuring no single party can tamper with the history of agent interactions or outcomes.
@@ 4.3.3 Secure Team Formation
An advanced feature of the Coral Protocol is its support for secure team formation, which is the dynamic assembly of multiple agents into a collaborative team to solve complex tasks. When Interaction Mediation and Task Management determine that a user’s request requires diverse expertise (for instance, a task needs one agent to gather data, another to analyze it, and another to verify the results), the protocol initiates secure team formation. This process involves selecting the appropriate set of Coralized AI Agents, establishing authenticated communication channels among them, and defining each agent’s role in the workflow. The Secure Multi-Agent Workflow and blockchain infrastructure play pivotal roles here: each agent’s identity and permissions can be verified against blockchain records or certificates, and any inter-agent contracts (such as payment splits or data access agreements) can be codified as smart contracts or logged transactions. The result is a trusted ad-hoc team of AI agents that can cooperatively work on the user’s task. Throughout the team’s operation, the Coral services enforce that information sharing is restricted to what is necessary for the task and that any interim results are reported back to the Task Management module. Once the collaborative task is complete, the team can be disbanded, with the outcome delivered to the user as a coherent result. Secure team formation thus enables scalability of problem-solving: the ecosystem can marshal multiple specialized agents together, while maintaining security and trust among agents that may have been contributed by different developers or organizations.
@@ 4.4   Inter-Context Data Flow and Interaction
Bringing together the above elements, the data flow in the Coral ecosystem moves fluidly across the three contexts under the governance of the Coral Protocol. A typical interaction begins with an end-user query or command from the Developers/Users context. This request enters the Coral Protocol, which uses its Interaction Mediation service to interpret and route the request to one or more Coralized AI Agents. Those agents, running on MCP Server infrastructure, receive the query along with any additional context or data provided (possibly accessing external data through integrated sources via Data Coralizers).
As the agents process the query, they may perform computations, look up information, or even generate sub-queries to other agents. Any such agent-to-agent interactions are again managed by the protocol’s workflow services. If the user’s request entails a transactional component (e.g. paying an agent or purchasing data), the Secure Payments service engages with the blockchain to execute and record the transaction, ensuring the agent proceeds with the task once payment is confirmed. For complex requests, the Task Management service may split the work among multiple agents and invoke secure team formation to coordinate them. Intermediate results flow back through the Task Management and Interaction Mediation layers, which assemble the final answer or result.
Finally, the response is delivered back to the end-user, completing the round-trip. Throughout this process, each component in the Developers/Users context (such as a developer’s integrated agent or data source) and each Coralized agent operates within the rules of the Coral Protocol, ensuring interoperability. The high-level architecture diagram (Figure 1) encapsulates these interactions, showing how the human-facing side, the AI agent side, and the protocol services side all connect. In essence, the Coral ecosystem’s architecture enables a seamless and secure loop: humans provide instructions and integrations, AI agents provide intelligence and action, and the Coral Protocol ties everything together with standardized communication, security, and coordination mechanisms suitable for a robust multi-agent system.
5   The Coral Protocol Architecture
@@ 5.1   Overview
The Coral Protocol’s architecture is organized into layers that connect end-user applications, developer tools, AI agents, and shared infrastructure. At a high level, user applications and developer tooling (top layer) interact with Coral Servers on each host, which in turn coordinate Coralized Agents (middle layer) running on distributed compute servers. These components communicate over the Internet and record trust-critical events on a common blockchain (bottom layer). Together, this architecture ensures that agents “wrapped” for Coral (the Coralized Agents) can be invoked and orchestrated in a uniform, secure manner.
Figure 2: The Coral Protocol Architecture
The architecture diagram of the Coral Protocol is shown in Figure 2. There are multiple computers (nodes) running Coral Servers, each hosting Coralized Agents, MCP servers, development tools, and wallets, all connected via the Internet and secured by an underlying blockchain.
@@ 5.2   Coralized Agents
The Coralized Agents reside in the middle tier of the architecture and are the core AI services of the system. Each Coralized Agent is an AI module or service (often backed by a large language model or specialized analytics tool) that has been onboarded to comply with Coral’s protocols (see Section 4.2). Once a model or data source is coralized, it becomes a first-class agent in the ecosystem. Collectively, these agents form a distributed, interoperable multi-agent system that can understand natural-language queries and perform tasks. When a user request or command arrives (via a user’s application), the Coral Server’s Interaction Mediation service dispatches the query to the appropriate Coralized Agent(s). For example, a query directed at “Agent X” is routed to that agent’s instance, which then processes the query and generates a response using its internal knowledge or tools. Coralized Agents can also communicate with one another (e.g. to handle subtasks in a workflow) under the coordination of the Coral Protocol’s multi-agent workflow services. In summary, the Coralized Agents layer encapsulates all AI services in the ecosystem, enabling them to interoperate seamlessly (see Section 4.2).
@@ 5.3   MCP Servers and Tools
The MCP Servers and Tools layer provides the compute and integration endpoints that Coralized Agents use to perform tasks. Coral leverages the Model Context Protocol (MCP) as a unified interface (see Section 2.4) to link agents with models and external tools. In practice, each agent is deployed as a service on one or more MCP servers. A network of MCP servers hosts AI model runtimes and heavy computation back-ends, exposing them through a standardized API. For instance, an MCP server might run a large language model or a data-processing pipeline; the Coral Protocol can invoke it remotely via the MCP interface. Notably, MCP supports not just model inference but also tool usage: agents can request to use external tools (APIs, databases, web search, etc.) through the same MCP API. In effect, “tools” in the diagram refers to additional functionalities or APIs that agents can call. By abstracting hardware and environment details, MCP servers ensure that any Coralized Agent can run anywhere in the network with proper context and resources. In short, the MCP Servers and Tools layer supplies the underlying AI models and function endpoints that agents use to generate intelligent results.
@@ 5.4   Coral Server
Internally, the Coral Server also invokes the Task Management service to decompose complex jobs into subtasks, and the Secure Multi-Agent Workflow to coordinate inter-agent protocols. If a transaction is involved, the Coral Server triggers the Secure Payments service to engage with the blockchain. In effect, the Coral Server is the control-plane nexus that ties user queries to agent actions: it routes inputs to the correct Coralized Agents, aggregates their outputs, and enforces security and workflow policies as described in Section 4.3.1.
@@ 5.5   Multi-Agent Application
The Multi-Agent Application layer sits atop the Coral Protocol on the developer’s computer. This is the developer’s or end-user’s application that uses Coral’s APIs to create and orchestrate agents. Typically, the application will instantiate one or more “UI agents” or orchestrator agents that interface with the human. For example, a UI agent can receive a user’s command (“Compute X for me”) and call the Coral Server to create a new task thread mentioning the required agents.

### v2

@@ 4   The Coral Ecosystem
The Coral ecosystem is organized into three major contexts: AI Agent Developers and Users, Coralized AI Agents, and Coral Protocol (see Figure 1). These contexts encapsulate the roles, components, and interactions that enable Coral’s secure and interoperable multi-agent environment. In this section, we describe each context and its components in detail, and explain the data flow and interactions between them in the overall architecture.
@@ 4.1   AI Agent Developers and Users
This context represents the human actors (developers and end-users) and their toolchain for interfacing with the Coral ecosystem. Developers contribute to Coral by integrating external AI capabilities and resources. They use specialized Coraliser modules to onboard different assets into the ecosystem: the MCP Coraliser connects external model endpoints via the Model Context Protocol (MCP), allowing external AI models or services to communicate in Coral’s standard format; the Data Coraliser links external data sources (e.g., databases, knowledge bases or live data streams) into the Coral ecosystem, making those data accessible to AI agents; and the Agent Coraliser wraps existing AI agents or services (such as pre-existing language models or automation scripts) to comply with Coral’s protocols and interfaces. Through these coralisers, developers effectively “Coralise” their models, data, or agents — registering them into the ecosystem so that they become available as Coralized AI Agents in the middle context. In addition to these, developers can utilize Other Tools (e.g., software development kits, testing frameworks, or monitoring dashboards provided by the Coral platform) to build, debug, and optimize their agents and integrations.
@@
End-users in this context are the consumers of the AI agent services. They interact with Coralized AI Agents by issuing natural language queries or commands through user-facing applications or interfaces built on top of Coral Protocol. Users need not be aware of the underlying complexity; they simply pose questions or tasks in everyday language. For instance, a user might ask, “What does AgentX think about hypothesisY?” or instruct the system with a command like, “Pay @AgentZ \$50 to investigate the total area of region R.” These inputs from users enter the Coral ecosystem via the Interaction Mediation service of Coral Protocol (described later), which ensures the queries are routed appropriately. In summary, the Developers and Users context supplies the ecosystem with integrated AI capabilities (via coralizers) and the driving queries or tasks (from users) that initiate agent interactions.
4.2   Coralised AI Agents
@@
The middle tier of the architecture consists of the Coralised AI Agents — the ensemble of AI services and agents that have been onboarded into the ecosystem and operate under Coral’s framework. Each Coralised agent is an AI service (often backed by a large language model or specialized AI module) that adheres to Coral Protocol for communication and security. Once developers coralise a model, data source, or existing agent, it becomes a Coralised AI Agent accessible in this layer. Collectively, these agents form a distributed, interoperable multi-agent system, each capable of understanding and responding to natural language prompts.
@@
When an end-user query or command enters the system (for example, the questions or instructions mentioned above), Coral Protocol’s Interaction Mediation component will dispatch it to the appropriate agent or agents in this layer. An agent specialized in a certain domain (say, Agent X in the earlier query) will receive the question directed at it and generate a response drawing on its knowledge or data (which might have been linked via a Data Coralizer). Coralized agents can also initiate interactions with one another under guidance of Coral Protocol. For instance, if a user’s request is complex and requires multiple skills, Coral Protocol can intelligently form a dedicated team by discovering and composing the necessary AI agent together. All such inter-agent dialogues are mediated and logged by the protocol to maintain coherence and security.
@@
Crucially, Coralized AI Agents can handle not only informational queries but also procedural commands that involve actions in the ecosystem. For example, when a user says “Pay @AgentZ \$50 to perform task T,” the addressed agent (Agent Z) will be invoked to execute the task T, and Coral Protocol will engage its secure payment service to transfer the specified amount. Throughout their operation, coralized agents remain decoupled from any particular user interface or platform — they rely on Coral Protocol to handle incoming requests, outgoing replies, and any coordination with other services. This design allows a heterogeneous collection of AI agents to work together seamlessly: each agent focuses on its specialized processing (e.g., analyzing data, answering questions, performing computations), while the protocol manages communications and shared context. The result is a flexible multi-agent ecosystem where, from the user’s perspective, complex tasks (potentially involving several agents and data sources) can be invoked with simple natural language requests.
@@ 4.3   Coral Protocol
Coral Protocol is the underlying framework that connects users, developers, and coralized agents, providing the services and infrastructure necessary for secure, coordinated interactions. It comprises two main parts: a set of core Coral Services that mediate and secure the ecosystem’s operations, and the Deployment and infrastructure layer that supports these services. An important additional concept in this context is Secure Team Formation, which refers to the protocol’s capability to dynamically assemble multiple agents into collaborative task teams in a secure manner. We describe each of these aspects below.
@@ 4.3.1 Coral Services
Coral Protocol offers several key services to orchestrate interactions:
@@
    Secure Multi-Agent Teamwork coordinates scenarios where multiple AI agents collaborate on a shared objective. This service sets up and manages the team of information and control between agents, enforcing security policies such as authentication of each agent’s identity and permissions. It ensures that agents exchange data through approved channels and that the execution follows any constraints (for instance, not revealing sensitive data to an agent lacking clearance). Together, these Coral services form the intelligent control plane of the ecosystem, mediating every interaction to guarantee it is executed efficiently and safely.
@@ 4.3.2 Deployment Infrastructure
To support the above services and the execution of AI agents, Coral Protocol leverages a distributed deployment infrastructure. A network of MCP Servers hosts the AI models and agent runtimes. “MCP” refers to the Model Context Protocol – a standardized interface that these servers implement to allow agents to request model inference or tool usage with a unified API. By deploying agents on MCP servers, Coral ensures that each agent can be invoked remotely with proper context and resource allocation, abstracting away the hardware or environment details. These servers handle the heavy computation of AI tasks (such as running large language model inferences or data processing jobs) and expose endpoints for the Interaction Mediation service to call. In parallel, the Coral ecosystem integrates a blockchain network into its infrastructure. The blockchain serves as a secure ledger for recording important events and transactions: it records payment transactions issued via the Secure Payments service, and can also log multi-agent agreements or task completion records for auditability. By using blockchain technology, Coral introduces an immutable and transparent layer of trust – for example, users and developers can verify that an agent was indeed paid for its service, or that a particular multi-agent task’s outcome was agreed upon by all parties. The blockchain also aids in decentralized governance of the ecosystem, ensuring no single party can tamper with the history of agent interactions or outcomes.
@@ 4.3.3 Secure Team Formation
An advanced feature of Coral Protocol is its support for secure team formation, which is the dynamic assembly of multiple agents into a collaborative team to solve complex tasks. When Interaction Mediation and Task Management determine that a user’s request requires diverse expertise (for instance, a task needs one agent to gather data, another to analyze it, and another to verify the results), the protocol initiates secure team formation. This process involves selecting the appropriate set of Coralized AI Agents, establishing authenticated communication channels among them, and defining each agent’s role in the complex task. The Secure Multi-Agent Team and blockchain infrastructure play pivotal roles here: each agent’s identity and permissions can be verified against blockchain records or certificates, and any inter-agent contracts (such as payment splits or data access agreements) can be codified as smart contracts or logged transactions. The result is a trusted ad-hoc team of AI agents that can cooperatively work on the user’s task. Throughout the team’s operation, the Coral services enforce that information sharing is restricted to what is necessary for the task and that any interim results are reported back to the Task Management module. Once the collaborative task is complete, the team can be disbanded, with the outcome delivered to the user as a coherent result. Secure team formation thus enables scalability of problem-solving: the ecosystem can marshal multiple specialized agents together, while maintaining security and trust among agents that may have been contributed by different developers or organizations.
@@ 4.4   Inter-Context Data Flow and Interaction
Bringing together the above elements, the data flow in the Coral ecosystem moves fluidly across the three contexts under the governance of Coral Protocol. A typical interaction begins with an end-user query or command from the Developers/Users context. This request enters Coral Protocol, which uses its Interaction Mediation service to interpret and route the request to one or more Coralized AI Agents. Those agents, running on MCP Server infrastructure, receive the query along with any additional context or data provided (possibly accessing external data through integrated sources via Data Coralizers).
As the agents process the query, they may perform computations, look up information, or even generate sub-queries to other agents. Any such agent-to-agent interactions are again managed by the protocol’s services. If the user’s request entails a transactional component (e.g., paying an agent or purchasing data), the Secure Payments service engages with the blockchain to execute and record the transaction, ensuring the agent proceeds with the task once payment is confirmed. For complex requests, the Task Management service may split the work among multiple agents and invoke secure team formation to coordinate them. Intermediate results flow back through the Task Management and Interaction Mediation layers, which assemble the final answer or result.
Finally, the response is delivered back to the end-user, completing the round-trip. Throughout this process, each component in the Developers/Users context (such as a developer’s integrated agent or data source) and each Coralized agent operates within the rules of Coral Protocol, ensuring interoperability. The high-level architecture diagram (Figure 1) encapsulates these interactions, showing how the human-facing side, the AI agent side, and the protocol services side all connect. In essence, the Coral ecosystem’s architecture enables a seamless and secure loop: humans provide instructions and integrations, AI agents provide intelligence and action, and Coral Protocol ties everything together with standardized communication, security, and coordination mechanisms suitable for a robust multi-agent system.
5   Coral Protocol Architecture
@@ 5.1   Overview
Coral Protocol’s architecture is organized into layers that connect end-user applications, developer tools, AI agents, and shared infrastructure. At a high level, user applications and developer tooling (top layer) interact with Coral Servers on each host, which in turn coordinate Coralized Agents (middle layer) running on distributed compute servers. These components communicate over the Internet and record trust-critical events on a common blockchain (bottom layer). Together, this architecture ensures that agents “wrapped” for Coral (the Coralized Agents) can be invoked and orchestrated in a uniform, secure manner.
Figure 2: Coral Protocol Architecture
The architecture diagram of Coral Protocol is shown in Figure 2. There are multiple computers (nodes) running Coral Servers, each hosting Coralized Agents, MCP servers, development tools, and wallets, all connected via the Internet and secured by an underlying blockchain.
@@ 5.2   Coralized Agents
The Coralised Agents reside in the middle tier of the architecture and are the core AI services of the system. Each Coralised Agent is an AI module or service (often backed by a large language model or specialized analytics tool) that has been onboarded to comply with Coral’s protocols (see Section 4.2). Once a model or data source is coralised, it becomes a first-class agent in the ecosystem. Collectively, these agents form a distributed, interoperable multi-agent system that can understand natural-language queries and perform tasks. When a user request or command arrives (via a user’s application), the Coral Server’s Interaction Mediation service dispatches the query to the appropriate Coralised Agent(s). For example, a query directed at “Agent X” is routed to that agent’s instance, which then processes the query and generates a response using its internal knowledge or tools. Coralised Agents can also communicate with one another (e.g., to handle sub-tasks in a complex task execution) under the coordination of Coral Protocol’s multi-agent teamwork services. In summary, the Coralised Agents layer encapsulates all AI services in the ecosystem, enabling them to interoperate seamlessly (see Section 4.2).
@@ 5.3   MCP Servers and Tools
The MCP Servers and Tools layer provides the compute and integration endpoints that Coralized Agents use to perform tasks. Coral leverages the Model Context Protocol (MCP) as a unified interface (see Section 2.4) to link agents with models and external tools. In practice, each agent is deployed as a service on one or more MCP servers. A network of MCP servers hosts AI model runtimes and heavy computation back-ends, exposing them through a standardized API. For instance, an MCP server might run a large language model or a data-processing pipeline; Coral Protocol can invoke it remotely via the MCP interface. Notably, MCP supports not just model inference but also tool usage: agents can request to use external tools (APIs, databases, web search, etc.) through the same MCP API. In effect, “tools” in the diagram refers to additional functionalities or APIs that agents can call. By abstracting hardware and environment details, MCP servers ensure that any Coralized Agent can run anywhere in the network with proper context and resources. In short, the MCP Servers and Tools layer supplies the underlying AI models and function endpoints that agents use to generate intelligent results.
@@ 5.4   Coral Server
Internally, the Coral Server also invokes the Task Management service to decompose complex jobs into sub-tasks, and the Secure Multi-Agent Teamwork to coordinate inter-agent protocols. If a transaction is involved, the Coral Server triggers the Secure Payments service to engage with the blockchain. In effect, the Coral Server is the control-plane nexus that ties user queries to agent actions: it routes inputs to the correct Coralised Agents, aggregates their outputs, and enforces security and workflow policies as described in Section 4.3.1.
@@ 5.5   Multi-Agent Application
The Multi-Agent Application layer sits atop Coral Protocol on the developer’s computer. This is the developer’s or end-user’s application that uses Coral’s APIs to create and orchestrate agents. Typically, the application will instantiate one or more “UI agents” or orchestrator agents that interface with the human. For example, a UI agent can receive a user’s command (“Compute X for me”) and call the Coral Server to create a new task thread mentioning the required agents.